OpenStudy (anonymous):

x+2 = RADICAL (7 + 2x)

OpenStudy (saifoo.khan):

(x+2)(x+2) = 7 +2x Sp;ve.

OpenStudy (saifoo.khan):

SOLVE*

OpenStudy (anonymous):

YES! Solve for x

OpenStudy (saifoo.khan):

x^2 + 4x + 4 -2x -7 = 0 x^2 + 2x -3 = 0

OpenStudy (anonymous):

(x+2)^2=7+2x ===> x^2+2x-3=0 ===> (x+3)(x-1)=0 ===> x--3 or x=1

OpenStudy (anonymous):

I get (x+1) (x-3), but when I plug them back in, nothing seems to work! HELP?

OpenStudy (anonymous):

I got x=-1 and x+3, but they don't seem to be solutions - please check them for me too! I'm losing my mind!

OpenStudy (saifoo.khan):

1, -3 Solutions.

OpenStudy (anonymous):

but, they don't work when you plug them back in to the original equation do they? I couldn't get them to work

OpenStudy (saifoo.khan):

yes they do.

OpenStudy (saifoo.khan):

try it.

OpenStudy (saifoo.khan):

3 = 3

OpenStudy (anonymous):

both of them?

OpenStudy (saifoo.khan):

the second one is messed up!

OpenStudy (saifoo.khan):

-1 = 1